OpenStudy (anonymous):

Would the difference of two infinite squares be indeterminate form? \[\sqrt{H + 1} - \sqrt{H + 2}\]

OpenStudy (anonymous):

\[ \infty - \infty \]Is an indeterminate form.

OpenStudy (anonymous):

\[\sqrt{\infty} \to \infty \]

OpenStudy (anonymous):

Infinity pulse constant is infinite? \[ \sqrt{\infty + b} = \sqrt{\infty} = \infty \]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What are you trying to do?